OpenStudy (bbb911):

Add or subtract. Then simplify by collecting like radical terms, if possible. Assume that all expressions under radicals represent nonnegative numbers.

OpenStudy (bbb911):

\[\sqrt{3} - 2\sqrt{108} + 5\sqrt{27}\]

rishavraj (rishavraj):

\[\it~simplifies~\sqrt{3} - 12\sqrt{3} + 15\sqrt{3}\]

OpenStudy (bbb911):

so you simplified now can it be more simplified ?

rishavraj (rishavraj):

\[\sqrt{3} = 1.732\]

OpenStudy (bbb911):

none of my answers have that answer

rishavraj (rishavraj):

\[its~4\sqrt{3}\]

OpenStudy (bbb911):

Thank you so much! :D

OpenStudy (bbb911):

I have one more question if you don't mind?

OpenStudy (bbb911):

ill post it

rishavraj (rishavraj):

no problem at all...

OpenStudy (bbb911):

yay thank you